EVA Report – January 6th

EVA #7

Author: Denys Bulikhov (EXO)

Purpose of EVA: Collection of geological samples and ambient radiation readings

Start time: 12:34

End time: 14:19

Narrative: EVA 7 crew took Cow Dung Road up to the intersection with Galileo Rd. The mountain on the left of the intersection was supposed to be the first stop. However, due to fresh snow the intersection was missed. EVA 7 crew had their first stop about 0.5 miles down the Cow Dung Rd past the intersection with the Galileo Rd. After collection of samples and ambient radiation readings, EVA 7 crew came back to the intersection with Galileo Rd. On the way back, EVA 7 crew stopped at what looked like a shallow river bed and collected samples. Final stop was made at the Marble Ritual.

Destination: intersection of Galileo and Cow Dung Rds.

Coordinates: N518900, E4252000

Participants: Denys Bulikhov (EXO), Alexandra Dukes (JOU), Jake Qiu (HSO)

Road(s) and routes per MDRS Map: Cow Dung Road.

Mode of travel: Driving and walking

Vehicles used: Spirit and Curiosity.

Sol Summary – January 5th

Crew 202 Sol Summary Report 05-Jan-2019

Sol: 7

Summary Title: Half-way

Author’s name: Cesare Guariniello

Mission Status: We are now half-way through our mission, and enjoying the warm habitat while expecting more snow overnight

Sol Activity Summary: To help recovering from the toughness of Mars, this morning we opted for yoga and guided meditation for our morning routine. Following the warmest EVA of the mission (temperatures still around freezing, but bright sun), today we had the coldest EVA of the mission, due to cloudy sky and cold wind. The EVA was intended to reach the areas of canyons at the end of Cactus Rd, but the snow made it difficult to recognize the trail in certain areas, so the EVA crew stopped along the way to work on their research project. After a hot lunch, the afternoon was spent watching a movie, working, and resting.

Look Ahead Plan: Tomorrow we will have a short EVA if precipitation ends and conditions are acceptable. Mostly, we will enjoy a day indoor.

Anomalies in work: None

Weather: Cloudy, expecting snow tonight and tomorrow morning

Crew Physical Status: Some tiredness here and there, but generally very good

EVA: The Commander, Executive Officer, Crew Geologist, and Crew Engineer had an EVA to the East, along Galileo Rd and the initial part of Cactus Rd. Condition of the trail did not allow to go all the way to the canyons region (snow made it very hard to discern the trail, and the Commander decided to stop in a region of interest along the way).

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, EVA report, EVA request, Astronomy Report, Journalist Report

Support Requested: None

Sol Summary – January 6th

Sol: 8

Summary Title: School starts tomorrow

Author’s name: Cesare Guariniello

Mission Status: Despite waking up to a white Mars again, the mission is proceeding very well, and we also managed to have a short EVA

Sol Activity Summary: Today, Sunday on Earth, the sol started a little slow. We woke up 30 minutes later than the usual schedule and had some great stretching. After sweeping the tunnels and enjoying some personal time, we had a short EVA to gather measurements on radiation. Two crew members took a shower today. The struggle with alternating icy cold and scalding hot water is real, but the aftermath is amazing! The tired crew enjoyed a relaxed afternoon, thinking of students at Purdue who will start their classes tomorrow

Look Ahead Plan: Tomorrow we will try again to have an EVA towards the canyon area

Anomalies in work: None

Weather: Sunny and warmer

Crew Physical Status: Two crew member are in perfect mood after the shower, four more crew members are doing good

EVA: The Executive Officer, Journalist, and Health and Safety Officer had an EVA along Galileo Rd and to Marble Ritual

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, EVA report, EVA request, Journalist Report

Support Requested: None